At 1:30pm an airplane leaves Tucson for Miami a distance of 2240 miles. The plane flies at 280 mph. A second airplane leaves Tucson at 2:15pm and is scheduled to land in Miami 15 minutes before the first airplane. At what rate must the second airplane travel to arrive on schedule? PLEASE HELP!!!! THANK-YOU
Answer by ptaylor(1894)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
amount of time it takes first plane to arrive in Miami=(d/r)=2240/280=8 hr.
This means that first plane gets into Miami at 9:30pm and we are told that the second plane is scheduled to land at 9:15pm. So, the second plane has to make the trip in (9:15-2:15)or 7 hrs. Now we can easily calculate the rate of the second plane:
Rate of second plane=(d/t)=2240/7=320 mph
Hope this helps---ptaylor